Which of the following pairs of functions are inverse functions?
a. f(x)=3x-9, g(x)=-3x+9
b. f(x)=4-x, g(x)=4+x
c. f(x)=x+5, g(x)=x-5
d. f(x)=x, g(x)=-x

Answer by Nate(3500)

c)
~
a*y = b*x + c with constants a, b, and c
Inverse: y = (a/b)x - (c/b)
~ example:
f(x) = 3x - 9
a = 1, b = 3, and c = -9
y = (1/3)x + (9/3)
y = (1/3)x + 3